# Spineless Traversal for Layout Invalidation

### v4

@@ 5.2. Dependent Synthesis
The most critical responsibility of the compiler is synthesizing code that sets dependent dirty bits every time a layout field is modified. This follows the standard approach from prior work. In short, for every field $U$, the compiler identifies all fields on all nodes that are affected when the value of $U$ changes, and generates code to set their dirty bits by calling the dirty method. To identify these affected fields, the compiler examines every other field assignment $\mathsf{self}.V\leftarrow T$ in the layout rules. For every field access $N.U$ in $T$, we know that $\mathsf{self}.V$ depends on $N.U$, meaning that any changes to $\mathsf{self}.U$ must mark $N^{-1}.V$ as dirty; here $N^{-1}$ is the inverse of the relation $N$; flipping next and previous, mapping first and last to parent, and mapping parent to all children.
Moreover, insertions or deletions must also set dirty bits. Inserting a node changes the meaning of the prev and next pointers for its siblings, and possibly the first and last pointers for its parent; deleting a node does the same. This means that inserting or deleting a node must dirty every field computation that reads any field from those pointers, and also any field that uses $\mathsf{N?}$ expressions. In Megatron, the analysis we perform is flow-insensitive; a real-world implementation might instead use a flow-sensitive analysis to dirty fewer nodes, but the engineering trade-offs are more challenging and we felt that the focus of this paper—the invalidation traversal—would have similar impact in either case.
@@
    The $H$ field depends on $\mathsf{last}?$, $\mathsf{last}.HA$, and $\mathsf{self}.\text{attribute}[\mathsf{height}]$, and the $\mathsf{last}?$ dependency is subsumed by the $\mathsf{last}.HA$ dependency. Taking inverses, we obtain: modifying $\mathsf{self}.\text{attribute}[\mathsf{height}]$ must dirty $\mathsf{self}.H$ and, if the node is a last child, modifying $\mathsf{self}.HA$ or inserting/deleting a node must dirty $\mathsf{parent}.H$.
@@
    The $HA$ field depends on $\mathsf{prev}?$, $\mathsf{prev}.HA$, and $\mathsf{self}.H$, with the $\mathsf{prev}$ dependency subsumed. Taking inverses, modifying $\mathsf{self}.H$ must dirty $\mathsf{self}.HA$101010 As described later, field packing would likely assign the $H$ and $HA$ fields the same dirty bit; in this case, modifying $\mathsf{self}.H$ need not dirty $\mathsf{self}.HA$, since it is already dirty. and, if the node is not a last child, modifying $\mathsf{self}.HA$ or inserting/deleting nodes must dirty $\mathsf{next}.HA$.
These modification rules, gathered from analyzing each field access, are grouped by which field is being modified and then injected into the relevant case of the recompute function. For example, the code to recompute $HA$ would check whether or not the node is a last child and dirty either $\mathsf{parent}.H$ (if so) or $\mathsf{self}.HA$ (if not). Additionally, Megatron makes sure to only dirty any given field once per field recomputation. For example, if a field $N.V$ is used twice in an expression, or if two different accesses $\mathsf{first}.V$ and $\mathsf{last}.V$ have the same inverse, the field is only dirtied once. This deduplication is especially challenging in the case of $N?$ expressions, since whether or not a field is dirtied can depend on whether the node is the first or last child of its parent. This deduplication is complex but, ultimately, possible to perform statically, which is critical to ensure maximal performance.
@@ 5.3. Code Generation
Megatron’s code generation has three key steps: generating the node data structure, the field value computations, and the top-level recompute function. To simplify code generation and optimization, the compiler is organized as a staged interpreter (MetaOCaml). That is, we first implemented an interpreter, and then added staging annotations that generate a custom IR; Figure 10 summarizes the approach. The custom IR is then converted to C++ using traditional compiler optimizations. This development approach was critical, as we found compiler correctness challenging: flaws in the dependency analysis could have unintuitive, long-distance consequences on program behavior that were difficult to distinguish from incorrect optimizations. The staging step allowed for much easier debugging.
@@
To generate efficient node data structures, Megatron type-checks all fields, attributes, and properties using Hindley-Milner type inference (Milner, 1978) and uses appropriate unboxed C++ member variables to store the relevant values. Importantly, this means a single node and all its fields are contiguous in memory (as it would be a real web browser) with a minimum of pointers (beyond the standard parent/first/last/next/previous pointers to other layout nodes) and that field access is compiled to a memory offset. All string values (like the keyword values for display) are interned and represented in C++ as a single enum type, meaning that no string allocation or comparison is performed at runtime. Discriminated unions are used for fields with units (for example $\text{property}[\mathsf{length}]$ can be an absolute length or a percentage). Dirty bits, summary bits, and timestamps are placed adjacent to the fields they cover.
To reduce overhead as much as possible, Megatron packs multiple co-computed fields, covering them with a single dirty bit. Packed fields are laid out adjacent in memory because they are written to, and often read from, in rapid succession. While more sophisticated techniques exist (Liu et al., 2023), Megatron uses a simple approach: each pass in the layout rules gets two dirty bits, one for the pre-order-computed fields and one for post-order-computed fields. The intuition is that this minimizeds the number of unique dirty bits set during dirty bit propagation and reduces the size of the priority queue and the number of order maintenance objects needed; such field packing is common in real browsers. Field packing has complex trade-offs not implemented in Megatron but Spineless Traversal would work with any field packing.
@@
6. Micro-Optimizations
Maximizing Spinless Traversal’s advantage over Double Dirty Bit requires careful micro-optimizations to improve cache locality, reduce memory traffic, and alleviate branch mispredictions. Unlike the optimization passes of the Megatron compiler, which apply to all invalidation algorithms, these optimizations are implemented in the Spineless Traversal invalidation algorithm and only affect it.
@@ 6.1. Pointer Compression and OM Allocation
Order maintenance objects have to be allocated every time new layout nodes are inserted; optimizing that allocation is essential. We use a hand-written pool allocator; in fact, we use separate pools for high- and low-level list cells to enhance locality. Since allocation are always the same size, and since layout is single-threaded, [^11] As it is in all major browsers. our custom allocator is significantly simpler than the system malloc.
The allocator, OMPool, is shown in Figure 11. It is parameterized by two types: the type of allocated object T and the index type P for pointers to allocated objects. Crucially, malloc and free return and consume the pointer type P instead of raw pointers. (And addressof function converts the pointer type P to a raw pointer so that it can be dereferenced.) By making P a small integer type, like uint32 or uint16, order maintenance objects become smaller and more of them fit per cache line, which in turn improves throughput. In our implementation, we use 32-bit pointers; this conveniently makes the total size of an order maintenance object 128 bits, meaning that no order maintenance objects split across two cache lines.
The actual implementation of OMPool is standard; it stores a pool of memory as a std::vector, in which we ensure sufficient capacity at startup. Freed elements are placed in a separate freed vector, which is preferentially drawn from by malloc. Because the objects are all the same size, there is no fragmentation and malloc/free are nearly instantaneous. Moreover, since Spineless Traversal creates Order Maintenance objects in order, allocation patterns are extremely favorable, with temporally-close nodes often placed nearby in memory.
template<  typename T,                   // Type of allocated object  typename P=uint32_t           // Integer type for "pointers"> struct OMPool {  std::vector<T> pool;          // Fast, local allocations  std::vector<P> freed;         // Rapid reuse minimizes churn  P malloc();                   // Straightforward  void free(P p) { freed.push_back(p); }  T* addressof(P p) { return &(poolp); }}
@@ 6.2. Branchless Order Maintenance Comparison
Priority queue pushes and pops spend basically all of their time comparing order maintenance objects. Order maintenance objects are small and, thanks to our allocator, typically cached. However, order maintenance object comparison has two cases (same or different second-level lists) and the bottleneck ends up being the pipeline stall induced by the conditional. The branch predictor does not help much, because (thanks to the priority queue) the comparison is unpredictable. We therefore implemented a branchless comparison function, relying on inline assembly cmov instructions, shown in Figure 12; it executes in about five cycles of latency. Assembly makes the implementation non-portable, but this function is critical for performance and we weren’t able to make the C++ compiler generate comparably-fast code, despite several attempts.
@@ 6.3. Attempted, Failed Optimizations
  • Pointer tagging to make priority queue elements smaller. No faster.
@@
  • Deallocating OM objects when the corresponding tree node is deleted, to increase cache locality. No improvement.
@@ 7. Evaluation
We implement a fragment of web layout in Megatron and use it to compare Spineless Traversal against the Double Dirty Bit algorithm on 50 real-world websites.
@@ 7.1. Web Layout Fragment
Existing web layout implementations are too complex and too tightly coupled to their current invalidation strategy, making them impossible to reuse. We instead implemented our own based on Cassius and Medea (Panchekha and Torlak, 2016; Panchekha et al., 2018; Liu et al., 2023). Naturally, our implementation handles only a subset of HTML and CSS features. However, we took care to implement several features with complex invalidation behavior, including intrinsic width, flexbox layout, and positioning. In total, our implementation computes approximately 50 layout fields over about 700 lines of Megatron DSL. This subsection describes several key layout features we implemented.
@@
Each layout node has an $x$, $y$, width and height field. Typically a node’s rectangle contains its children and doesn’t overlap with siblings. Width is generally has parent-to-child dependencies, height has child-to-parent dependencies, and $x$ and $y$ are computed in-order. This forms long dependency chains between elements—modifying one element can eventually dirty many others—but many CSS properties like width, min-width, and max-width, and similar for height, can break these dependency chains. These properties, however, allow values like 50%, which are resolved relative to the parent and thus still creates inter-node dependencies.
@@
Line breaking lays out inline layout nodes (text) horizontally into lines, until the edge of the parent box is reached, and which point the next inline layout node is placed in the next line. Line breaking thus creates control dependencies, where the parent node’s width may move layout nodes from one line to another by changing line breaking. Additionally, our layout algorithm allows different lines to have different heights (based on the height of the largest layout node in the line), which introduces a field (line height) that is dependent on many different nodes (each word in the line). This has a number of interesting effects for invalidation. For example, adding a node to a line may or may not change the line’s height, depending on whether the new node is tallest. If it is, all other text on the page must move down, causing a lot of invalidation.
@@
An element with position: absolute is manually assigned its $x$ and $y$ position by the web developer; this property is used for popups, tool-tips, and other hover effects. It is also common to change the manually-assigned $x$ and $y$ positions from JavaScript, such as to move a tool-tip away from the cursor. Layout nodes with absolute positioning do not affect the position of other layout nodes outside their subtree, and handling changes to $x$ and $y$ positions quickly is essential.
@@
Layout node have an “intrinsic” size—its size without line breaks, basically—which is used for, for example, absolutely-positioned elements. Importantly, intrinsic widths are computed bottom-up, but is then used in the top-down width computation, which then affects the bottom-up height computation. This means intrinsic sizes require the use of multiple layout phases.
@@
Flexbox layout is the most complex feature we implemented. In flexbox layout there is flex container element whose children are flex items. The width/height of flex items depends on the intrinsic sizes of the other flex items and the actual size of the flex container. Properties like flex-grow and flex-shrink determine how the intrinsic sizes of the flex items are adjusted to compensate for either extra or not enough available space in the flex container. The max- and min-width/height properties can also cap the growth/shrinkage of individual flex items. In all, our implementation of flexbox layout uses 9 intermediate fields and requires 2 passes to compute all of them.
@@
We also implemented a variety of miscellaneous features, including automatic sizing of images and video, manual line breaks with the ¡br¿ element, and hidden elements like ¡noscript¿ (which are only rendered if JavaScript is disabled). We also had to add a special case for ¡svg¿ elements, whose children describe drawing commands that do not participate in layout. Finally, we also implemented the width and height HTML attributes (which behave slightly differently from the CSS properties).
7.2. Benchmarks
@@ 7.3. Results
Figures 14 and 14 includes both “overhead” and “evaluation” time. Breaking those out further, with Spineless Traversal, overhead time was roughly one third of total runtime, while evaluation time was roughly two thirds, showing that invalidation overhead is still a significant determinant of latency. Naturally, the slower Double Dirty Bit algorithm spends even more time in invalidation overhead. Evaluation time, as expected, is nearly identical between Spinless Traversal and Double Dirty Bit. Meanwhile, considering overhead alone, Spineless Traversal as, for some frames, as much as 100×100$\times$100 × faster than Double Dirty Bit.
A more careful inspection of Figure 14 shows several additional features. The slowest frames all feature slowdowns. This is expected: the slowest frames likely represent the initial page load or other “loading” frames, which we necessarily capture in our traces. While speedups are always better than slow-downs, these loading frames likely follow network latency, so invalidation time for these frames is less important. Meanwhile, frames where fewer nodes are invalidated are typically those triggered in response to an animation or user interaction, where latency is most critical. When we restrict the data to only include frames where fewer than 1% of fields are recomputed—the intention being to ignore “loading” frames—the geometric mean speedup is larger, at 2.35×2.35$\times$2.35 ×, and a smaller fraction of frames suffer slowdowns. Spineless Traversal is also faster outside this subset, likely because the 1% threshold is an imprecise heuristic, but the larger speedup in this latency-critical subset is indicative.

### v5

@@ 5.2. Dependent Synthesis
The most critical responsibility of the compiler is synthesizing code to sets dependent dirty bits every time a layout field is modified. Megatron follows the standard approach from prior work. In short, for every field $U$, the compiler identifies all fields on all nodes that are affected when the value of $U$ changes, and generates code to set their dirty bits by calling the dirty method. To identify affected fields, the compiler examines every other field assignment $\mathsf{self}.V\leftarrow T$ in the layout rules. For every field access $N.U$ in $T$, we know that $\mathsf{self}.V$ depends on $N.U$, meaning that any changes to $\mathsf{self}.U$ must mark $N^{-1}.V$ as dirty; here $N^{-1}$ inverts the relation $N$ by flipping next and previous, mapping first and last to parent, and mapping parent to all children.
Insertions and deletions also set dirty bits. Inserting a node changes the meaning of the prev and next pointers for its siblings, and possibly the first and last pointers for its parent; deleting a node does the same. This means that inserting or deleting a node must dirty every field computation that reads any field from those pointers, and also any field that uses $\mathsf{N?}$ expressions. In Megatron, the analysis we perform is flow-insensitive; a real-world implementation might instead use a flow-sensitive analysis to dirty fewer nodes, but the engineering trade-offs are more challenging and we felt that the focus of this paper—the invalidation traversal—would have similar impact in either case.
@@
    The $H$ field depends on $\mathsf{last}?$, $\mathsf{last}.HA$, and $\mathsf{self}.\text{attribute}[\mathsf{height}]$, and the $\mathsf{last}?$ dependency is subsumed by the $\mathsf{last}.HA$ dependency. Taking inverses, modifying $\mathsf{self}.\text{attribute}[\mathsf{height}]$ must dirty $\mathsf{self}.H$ and, if the node is a last child, modifying $\mathsf{self}.HA$ or inserting/deleting a node must dirty $\mathsf{parent}.H$.
@@
    The $HA$ field depends on $\mathsf{prev}?$, $\mathsf{prev}.HA$, and $\mathsf{self}.H$, with the $\mathsf{prev}?$ dependency subsumed. Taking inverses, modifying $\mathsf{self}.H$ must dirty $\mathsf{self}.HA$101010 As described later, field packing would likely assign the $H$ and $HA$ fields the same dirty bit; in this case, modifying $\mathsf{self}.H$ need not dirty $\mathsf{self}.HA$, since it is already dirty. and, if the node is not a last child, modifying $\mathsf{self}.HA$ or inserting/deleting nodes must dirty $\mathsf{next}.HA$.
These modification rules, gathered from analyzing each field access, are grouped by which field is being modified and then injected into the relevant case of the recompute function. For example, the code to recompute $HA$ would check whether or not the node is a last child and dirty either $\mathsf{parent}.H$ (if so) or $\mathsf{next}.HA$ (if not). Additionally, Megatron makes sure to only dirty any given field once per field recomputation. For example, if a field $N.V$ is used twice in an expression, or if two different accesses $\mathsf{first}.V$ and $\mathsf{last}.V$ have the same inverse, the field is only dirtied once. This deduplication is especially challenging in the case of $N?$ expressions, since whether or not a field is dirtied can depend on whether the node is the first or last child of its parent. This deduplication is complex but, ultimately, possible to perform statically, which is critical to ensure maximal performance.
@@ 5.3. Code Generation
Megatron’s code generation has three key steps: generating the node data structure, generating the field value computations, and generating the top-level recompute function. To simplify code generation and optimization, the compiler is organized as a staged interpreter (MetaOCaml). That is, we first implemented an interpreter, and then added staging annotations that generate a custom IR; Figure 10 summarizes the approach. The custom IR is then converted to C++ using traditional compiler optimizations. This approach was critical, as we found compiler correctness challenging: flaws in the dependency analysis could have unintuitive, long-distance consequences on program behavior that were difficult to distinguish from incorrect optimizations. The staging step allowed for much easier debugging.
@@
To generate efficient node data structures, Megatron type-checks all fields, attributes, and properties using Hindley-Milner type inference (Milner, 1978) and uses appropriate unboxed C++ member variables to store the relevant values. Importantly, this means a single node and all its fields are contiguous in memory (as it would be a real web browser) with a minimum of pointers (beyond the standard parent/first/last/next/previous pointers to other layout nodes), with field access compiled to a memory offset. All string values (like the keyword values for display) are interned and represented in C++ as a single enum type, meaning that no string allocation or comparison is performed at runtime. Discriminated unions are used for fields with units (for example $\text{property}[\mathsf{length}]$ can be an absolute length or a percentage). Dirty bits, summary bits, and timestamps are placed adjacent to the fields they cover.
To reduce overhead as much as possible, Megatron packs multiple co-computed fields, covering them with a single dirty bit. Packed fields are laid out adjacent in memory because they are written to, and often read from, in rapid succession. While more sophisticated techniques exist (Liu et al., 2023), Megatron simply uses two dirty bits for each in the layout rules: one for the pre-order-computed fields and one for post-order-computed fields. The intuition is that this minimizes the number of unique dirty bits set during dirty bit propagation and reduces the size of the priority queue and the number of order maintenance objects needed. Field packing has complex trade-offs not implemented in Megatron but common in real browsers; however, Spineless Traversal would work with any field packing.
@@
6. Optimizations
Maximizing Spinless Traversal’s advantage over Double Dirty Bit requires careful optimizations to reduce work needed, improve cache locality, reduce memory traffic, and alleviate branch mispredictions. Unlike the optimization passes of the Megatron compiler, which apply to all invalidation algorithms, these optimizations are implemented in the Spineless Traversal invalidation algorithm and only affect it.
6.1. Queue Compression
Although the queue size is typically small, often only tens of elements long, it can occasionally become very large. This typically happens when the entire page must be re-laid-out, such as when the browser window is resized. In this case, every dirty bit ends up in the priority queue, and the queue can grow to thousands of elements long. This not only means that more queue pushes and pops are needed, but also that those pushes and pops take longer.
In theses cases the queue is often dense: for every dirty bit in the queue, its successor is also likely dirty and enqueued as well. Leveraging this insight, we can shrink the queue by representing sequences of enqueued dirty bits implicitly. That is, every time a queue element is popped, Megatron not only recomputes those fields but it also checks if the succeeding dirty bit is dirty and, if so, recompute it too. Megatron continues recomputing successive dirty bits until it reaches a cleared dirty bit. When a dirty bit is set, it doesn’t need to be enqueued if its predecessor is dirty. This leads to a significant reduction in queue length, and thus speeds up cases where many fields have to be recomputed, such as during initial loads and bulk inserts.
In principle, Megatron could even avoid checking predecessor dirty bits when this optimization is statically known to apply. However, we did not implement this further optimization.
@@ 6.2. Pointer Compression and OM Allocation
Order maintenance objects have to be allocated every time new layout nodes are inserted; optimizing that allocation is essential. We use a hand-written pool allocator; in fact, we use separate pools for high- and low-level list cells to enhance locality. Since allocations are always the same size, and since layout is single-threaded, [^11] As it is in all major browsers. our custom allocator is significantly simpler than the system malloc.
The allocator, OMPool, is shown in Figure 11. It is parameterized by two types: the type of allocated object T and the index type P for pointers to allocated objects. Crucially, malloc and free return and consume the pointer type P instead of raw pointers. (The addressof function converts the pointer type P to a raw pointer so that it can be dereferenced.) Making P a small integer type, like uint32 or uint16, shrinks order maintenance objects, which allows more of them fit per cache line, improving throughput. In our implementation, we use 32-bit pointers; this conveniently makes the total size of an order maintenance object 128 bits, meaning that order maintenance objects do not split across cache lines.
The actual implementation of OMPool is standard; it stores a pool of memory as an std::vector, in which we ensure sufficient capacity at startup. Freed elements are placed in a separate freed vector, which is preferentially drawn from by malloc. Because the objects are all the same size, there is no fragmentation and malloc/free are nearly instantaneous. Moreover, since Spineless Traversal creates Order Maintenance objects in order, allocation patterns are extremely favorable, with temporally-close nodes often placed nearby in memory.
template<  typename T,                   // Type of allocated object  typename P=uint32_t           // Integer type for "pointers"> struct OMPool {  std::vector<T> pool;          // Fast, local allocations  std::vector<P> freed;         // Rapid reuse minimizes churn  P malloc();                   // Straightforward  void free(P p) { freed.push_back(p); }  T* addressof(P p) { return &(pool[p]); }}
@@ 6.3. Branchless Order Maintenance Comparison
Priority queue pushes and pops spend basically all of their time comparing order maintenance objects. Order maintenance objects are small and, thanks to our allocator, typically in cache. However, order maintenance object comparison has two cases (same or different second-level lists) and the pipeline stall from the conditional ends up being a bottleneck. Moreover, the branch predictor does not help much, because (thanks to the priority queue) the comparison is unpredictable. We therefore implemented a branchless comparison function, relying on inline assembly cmov instructions, shown in Figure 12; it executes in about five cycles of latency. Assembly makes the implementation non-portable, but this function is critical for performance and we weren’t able to make the C++ compiler generate comparably-fast code, despite several attempts.
@@ 6.4. Attempted, Failed Optimizations
  • Pointer tagging to make priority queue elements smaller. No improvement.
@@
  • Deallocating OM objects when the corresponding tree node is deleted, to increase cache locality. A slight slow-down.
@@ 7. Evaluation
We implement a fragment of web layout in Megatron and use it to compare Spineless Traversal against Double Dirty Bit on 50 real-world websites.
@@ 7.1. Web Layout Fragment
Existing web layout implementations are complex and tightly coupled to their current invalidation strategy. Therefore, to evaluate Spineless Traversal, we re-implemented web layout, basing our approach on Cassius and Medea (Panchekha and Torlak, 2016; Panchekha et al., 2018; Liu et al., 2023). Naturally, our implementation handles only a subset of HTML and CSS features. However, we took care to implement several features with complex invalidation behavior described below. In total, our implementation computes approximately 50 layout fields over about 700 lines of Megatron DSL.
@@
Each layout node has $x$, $y$, width and height fields; formally, this rectangle defines its border box. Typically a node’s border box contains its children and doesn’t overlap with siblings. Width generally has parent-to-child dependencies, height generally has child-to-parent dependencies, while $x$ and $y$ are computed in-order. This forms long dependency chains between elements—modifying one element can eventually dirty many others—but many CSS properties like width, min-width, and max-width, and similar for height, can break these dependency chains. These properties, however, allow values like 50%, which are resolved relative to the parent and thus still creates inter-node dependencies.
@@
Line breaking lays out inline layout nodes (text) horizontally into lines. When text reaches the right edge of its parent box, the next inline layout node is placed in the next line. Line breaking thus creates control dependencies, where checking the parent node’s width may cause layout nodes to move from one line to another (by changing line breaking). Additionally, our layout algorithm allows different lines to have different heights (based on the height of the largest layout node in the line), which introduces a field (line height) that is dependent on many different nodes (each word in the line). This also requires multiple layout passes, since later words in a line can affect the placement of earlier words by adjusting the text baseline. This has a number of interesting effects for invalidation. For example, adding a node to a line may or may not change the line’s height, depending on whether the new node is tallest. If it is, all other text on the page must move down, causing a lot of invalidation.
@@
An element with position: absolute is manually assigned its $x$ and $y$ position by the web developer; this property is used for popups, tool-tips, and other hover effects. It is also common to change the manually-assigned $x$ and $y$ positions from JavaScript, such as to move a tool-tip away from the cursor. Layout nodes with absolute positioning do not affect the position of sibling layout nodes, and handling changes to $x$ and $y$ positions quickly is essential.
@@
Layout nodes have an “intrinsic” size—its size without any or with all line breaks, basically—which is used for, for example, absolutely-positioned elements. Importantly, intrinsic widths are computed bottom-up, but are then used in the top-down width computation, which then affects the bottom-up height computation. This means intrinsic sizes require the use of multiple layout phases.
@@
Flexbox layout is the most complex feature we implemented. In flexbox layout there is flex container element whose children are flex items. The width/height of flex items depends on the intrinsic sizes of the other flex items and the actual size of the flex container. Properties like flex-grow and flex-shrink determine how the intrinsic sizes of the flex items are adjusted to match available space in the flex container. The max- and min-width/height properties can also cap the growth/shrinkage of individual flex items. In all, our implementation of flexbox layout uses 9 intermediate fields and requires 2 passes to compute all of them. Note that the full web layout specification includes layout modes like grid layout, which require even more passes.
@@
We also implemented a variety of miscellaneous features, including automatic sizing of images and video, manual line breaks with the <br> element, and hidden elements like <noscript> (which are only rendered if JavaScript is disabled). We also had to add a special case for <svg> elements, whose children describe drawing commands that do not participate in layout. Finally, we also implemented the width and height HTML attributes (which behave slightly differently from the CSS properties).
7.2. Benchmark Web Pages
@@ 7.3. Results
Figures 14 and 14 includes both “overhead” and “evaluation” time. Evaluation time, as expected, is nearly identical between Spinless Traversal and Double Dirty Bit. Considering overhead alone, Spineless Traversal is, for some frames, as much as 100×100$\times$100 × faster than Double Dirty Bit. For Spineless Traversal, overhead time was roughly one third of total runtime, while evaluation time was roughly two thirds, showing that invalidation overhead is still a significant determinant of latency. Naturally, the slower Double Dirty Bit algorithm spends even more time in invalidation overhead. Breaking overhead time down further for Spineless Traversal shows that both the priority queue and the order maintenance structure contribute to overhead, with different algorithms dominating for different benchmarks
A careful inspection of Figure 14 shows several additional features. The slowest frames all feature slowdowns. This is expected: the slowest frames likely represent the initial page load or other “loading” frames, which we necessarily capture in our traces. While speedups are always better than slow-downs, these loading frames likely follow network latency, so invalidation time for these frames is less important. Meanwhile, frames where fewer nodes are invalidated are typically those triggered in response to an animation or user interaction, where latency is most noticeable as “jank”. When we restrict the data to only include frames where fewer than 1% of fields are recomputed—the intention being to ignore “loading” frames—the geometric mean speedup is larger, at $2\times2$, and a smaller fraction of frames (10%) suffer slowdowns. Spineless Traversal is also faster outside this subset, likely because the 1% threshold is an imprecise heuristic, but the larger speedup in this latency-critical subset is indicative.